# Game-Theoretic Coordination For Time-Critical Missions of UAV Systems

### v1

@@ Abstract
Cooperative missions involving Unmanned Aerial Vehicles (UAVs) in dynamic environments pose significant challenges in ensuring both coordination and agility. In this paper, we introduce a novel game-theoretic approach for time-critical missions, where each UAV optimizes a cost function that incorporates temporal and mission-specific constraints. The optimization is performed within a one-dimensional domain, significantly reducing the computational cost and enabling real-time application to complex and dynamic scenarios. The framework is distributed in structure, allowing to achieve global, system-wide coordination (a Nash equilibrium) by using only local information. For ideal systems, we prove the existence and exponential stability of the Nash equilibrium. Furthermore, we invoke model predictive control (MPC) for non-ideal scenarios. In particular, we propose a discrete-time optimization approach that tackles path-following errors and communication failures, ensuring reliable and agile performance in dynamic and uncertain environments. Simulation results demonstrate the effectiveness and agility of the approach in ensuring successful mission execution across diverse scenarios. Experiments using a motion capture system provide further validation under realistic conditions.
@@ I Introduction
Recent advancements in technology have significantly expanded the capabilities of Unmanned Aerial Vehicle (UAV) systems, enabling them to perform a wide range of tasks [27, 35]. Particularly, cooperative UAV systems have been extensively deployed across diverse domains, including civilian applications such as surveillance, environmental monitoring, and air traffic management, as well as military operations like swarm-based attack-defense scenarios. Notable examples include cooperative forest fire monitoring and suppression [28, 34], multiple moving targets surveillance [14], to name a few. Despite their broad range of applications, cooperative UAV systems face operational limitations and challenges like battery endurance, payload carrying capability, flight autonomy, path planning, path following, and achieving reliable cooperation.
Figure 1: Snapshot of the flight experiment demonstrating four UAVs coordinating and avoiding collisions in the presence of non-ideal communication. The added trail effect highlights the flight path of UAVs, a circular symbol representing the balance between two opposing forces, yin and yang.
@@
Effective cooperation among a system of UAVs is crucial for accomplishing complex tasks that exceed the capabilities of a single UAV. Applications such as simultaneous target tracking [3], formation flying [23], and large-scale area mapping and monitoring [32] heavily rely on the coordinated efforts of multiple UAVs. Achieving reliable cooperation, however, is inherently challenging due to factors such as communication delays, limited computational resources, variations in UAV dynamics, disturbances, and unpredictable events. These challenges highlight the need for efficient and adaptive strategies to ensure system-wide collaboration in dynamic and uncertain environments. This paper addresses the challenge of time-critical cooperation through a novel game-theoretic method.
Inspired by the pioneering work of Leslie Lamport on time synchronization [21], we formulate the time-critical coordination problem as a “time synchronization” problem (see Section II for more details). More precisely, following the framework of cooperative path-following in [2, 15, 18, 9], we decouple the space and time in the general problem formulation, which reduces the coordination problem to a one-dimensional consensus problem, thereby significantly lowering the computation costs. In [15, 17, 8, 18, 19], exponential stability of time coordination was established with a PI controller for networks connected in an integral sense. In this paper, we formulate the coordination problem as a game, and we use the virtual time (see [18]) as a consensus parameter for UAVs that engage in a game to reach an agreement on this parameter. As compared to the PI controller used in earlier work for coordination, we now provide a general approach to handle operational requirements and time-varying objectives. Furthermore, certain constraints, such as UAV speed limits or acceleration limits, may be directly incorporated into the optimization problem to ensure the feasibility of the provided solution. The optimization problem also enables the direct incorporation of complex, time-varying mission specifications into the optimization framework (e.g., loss of energy efficiency that might require mission modification or accommodating new collision avoidance requirements not accounted for in the apriori mission planning, as demonstrated in Section III-B, Remark 2). Figure 2 shows the architecture of the solution. The main difference from prior papers [2, 15, 18], which used a PI control law for coordination without considering optimality, lies in the block on Optimal Temporal Coordination, where a game theoretic formulation is considered for the UAVs to reach agreement on the consensus variables. The existence of a Nash equilibrium is shown and its exponential stability is proven.
Next, we highlight the main contributions of this paper:
  • We propose a novel game-theoretic coordination algorithm, where the Nash equilibrium of the strategic interactions among agents ensures system-wide coordination and successful task execution.
  • We prove the existence of a Nash equilibrium and establish the exponential stability under ideal conditions.
  • Through extensive simulations and flight experiments, we demonstrate the effectiveness of the algorithm for various coordination tasks in UAV systems.
In contrast to the classical Nash equilibrium existence results (see, for example, [24] and [37]), we prove the existence of continuous-time, infinite horizon Nash equilibrium in infinite-dimensional action domain with constraints. A similar result, like the existence of a Nash equilibrium with bounded control inputs and stability of the solution, was established in [36] using Lyapunov stability techniques. In this paper, we prove the existence and exponential stability of a Nash equilibrium with constraints, relying on Euler-Lagrange equations. For the Nash equilibrium analysis, Euler-Lagrange equations were used in [10] in the case of a single integrator. Here, we consider double integrator, where the corresponding system of Euler-Lagrange equations has a higher degree, and we study this system by explicitly solving it. Regarding Nash equilibrium seeking algorithms, most of the known results assume a fixed communication network (see, for example [25, 33, 13]). In contrast, leveraging MPC, we propose an algorithm that converges to Nash equilibrium even in the case of time-varying network.
@@
The remainder of this paper is structured as follows: Section II introduces the necessary notations and provides an overview of the preliminaries, including path planning, path following and coordination. Furthermore, in Section II, we introduce an infinite horizon game problem with discount factor. In Section III, by leveraging the system of Euler-Lagrange equations (a coupled fourth-order differential equations), we prove the existence of a Nash equilibrium for the corresponding unconstrained problem under ideal conditions. Moreover, we derive an explicit expression for the Nash equilibrium and prove the exponential stability of the solution; see Proposition 2. Eventfully, examining the coefficients of the explicit solution, we prove that there exists a discount rate such that the solution to the unconstrained problem is also solution to the constrained one; see Theorem 3. Section III-B outlines the general algorithm derived from the theoretical results, and Section IV validates the algorithm’s reliability and performance through simulations in challenging and realistic scenarios. Finally, Section V describes the flight experiments conducted with small UAVs. The video footage showcasing the simulations and the flight experiments is provided in the supplementary materials.
@@ II Preliminaries
We consider a class of time-critical cooperative missions that can be formulated as “time synchronization” problems (see, for example, [21, 18]). We assume that for a given mission a path planner generates a desired and feasible trajectory (possibly optimal for the mission) for each UAV. Next, we assume that each UAV is equipped with a robust low-level controller for tracking purposes. The coordination task is then reduced to a “time synchronization” problem.
Figure 2: Game-theoretic cooperative path-following of a system of UAVs: the block on Optimal Temporal Coordination uses a game-theoretic formulation and ensures existence and convergence of Nash equilibrium.
@@ II-A Path Planning and Following
Consider a system of $N$ UAVs involved in the mission, $N\in\mathbb{N}$. For each UAV the desired trajectory is given as a function $x_{d,i}:[0,t_{d,i}^{*}]\rightarrow\mathbb{R}^{3}$, $t_{d,i}^{*}>0$. For all $t\in[0,t_{d,i}^{*}]$, the trajectories satisfy the constraints defined by the $i^{\rm th}$ UAV’s dynamics:
$0\leqslant v^{i}_{\min}<v^{i}_{d,\min}\leqslant\left|\left|\dot{x}_{d,i}(t)\right|\right|\leqslant v^{i}_{d,\max}<v^{i}_{\max},$ (1)
@@
$\left|\left|\ddot{x}_{d,i}(t)\right|\right|\leqslant a^{i}_{d,\max}<a^{i}_{\max},$ (2)
@@
Time-critical mission specifications are encoded into the desired trajectories of the UAVs by the planner. By properly defining these trajectories, it is possible to design missions with varying characteristics, such as simultaneous arrival and/or sequential autolanding. The analysis of such systems are similar, and hence, here, we consider only the simultaneous arrival case. A simultaneous arrival mission is defined as:
$t_{d,i}^{*}=t_{d}^{*},\quad i=1,\dots,N.$
The trajectories must maintain spatial separation throughout the mission:
$\min_{\begin{subarray}{c}i,j=1,\ldots,n\\ i\neq j\end{subarray}}\left\|x_{d,i}(t)-x_{d,j}(t)\right\|^{2}\geqslant E^{2}>0,\quad\text{for all }t\geqslant 0.$ (3)
Remark 1. In Section III-B, we introduce an additional collision avoidance term, which allows to remove the spatial separation constraints in (3).
@@ II-B Time Coordination for Simultaneous Arrival
To formulate the problem of simultaneous arrival of UAVs, we follow [8, 18]. Towards that end, let $\gamma_{i}:\mathbb{R}^{+}\rightarrow[0,t_{d}^{*}]$, $i=1,\dots,N$, map the actual (clock) time $t$ to the planned mission time $t_{d}$, that is, $\gamma_{i}(t)=t_{d}$. In [8, 18], $\gamma_{i}$ is referred to as virtual time, which allows to redefine the desired trajectory of the $i^{\rm th}$ UAV as:
$x_{\gamma,i}\left(t\right)=x_{d,i}\left(\gamma_{i}(t)\right).$ (4)
Note that by this reparametrization of the desired trajectories, we gain control over the speed of each UAV, and hence, its mission execution time. Specifically,
  • When $\dot{\gamma}_{i}(t)=1$, the UAV travels at the desired pace.
  • When $\dot{\gamma}_{i}(t)>1$, the UAV moves faster than the desired pace.
  • When $\dot{\gamma}_{i}(t)<1$, the UAV moves slower than the desired pace.
Thus, $\gamma_{i}$ serves the role of a consensus parameter in our problem. Next, based on the physical limitations and mission requirements of the UAVs, as defined in (1) and (2), we derive general bounds for the consensus parameter of each UAV. Specifically, from (4), we have
@@
The first equation along with (1) implies that the minimum and the maximum values of all admissible parameters $\gamma_{i}$ must satisfy the following constraint:
@@
Note that since $v^{i}_{d,\max}<v^{i}_{\max}$ and $a^{i}_{d,\max}<a^{i}_{\max}$, there exist $\gamma_{i}$ parameters such that inequalities in (6) and (7) are satisfied.
@@ II-C Problem Formulation
Next, we introduce a game-theoretic approach that facilitates time-critical cooperative mission execution. For the theoretical analysis, here, we consider the path-following kinematics (ideal path-following) and a completely connected communication network. While these assumptions simplify the theoretical analysis (see Section III-A), they do not limit the practical applicability of the proposed approach. In Section III-B, we introduce an algorithm that incorporates path-following errors and communication failures, enabling application in real-world scenarios. As shown in [7, 16], by using cascaded inner-loop outer-loop structure for flight control applications, the uncertainties in system dynamics can be handled by a variety of robust control methods.
We begin by introducing the admissible set of the consensus parameter (virtual time) for each UAV. We have $\gamma_{i}(0)=\gamma^{0}_{i}\geqslant 0$, for $i=1,\dots,N$. Along with the inequalities in (6), (7), we define the admissible sets:
$\mathcal{A}^{0}_{i}:=\Big{\{}\gamma_{i}\in H^{1}_{w,\alpha}((0,\infty)):\gamma_{i}(0)=\gamma^{0}_{i}\geqslant 0,\,\dot{\gamma}_{i}(0)=1\Big{\}},$ (10)
@@
where $H^{1}_{w,\alpha}((0,\infty))$ is weighted Sobolev space (for more details see Appendix VIII) with weight $e^{-\alpha t}$
$\begin{split}H^{1}_{w,\alpha}((0,\infty)):=\Bigg{\{}g\in H^{1}_{loc}((0,\infty)):\\ \int_{0}^{\infty}e^{-\alpha t}(g^{2}+\dot{g}^{2}){\rm d}t<\infty\Bigg{\}}.\end{split}$ (11)
@@
$\begin{split}J_{i}(\gamma_{i},\gamma_{-i})=\int_{0}^{\infty}e^{-\alpha t}\Bigg{(}\left(\dot{\gamma}_{i}-1\right)^{2}+\sum_{j=1}^{N}(\gamma_{i}-\gamma_{j})^{2}+\ddot{\gamma}_{i}^{2}\Bigg{)}{\rm d}t,\end{split}$ (12)
where $\gamma_{-i}=(\gamma_{1},\dots,\gamma_{i-1},\gamma_{i+1},\dots,\gamma_{N})$.
The cost function in (12) corresponds to simultaneous arrival and maintaining a desired mission pace with the following terms:
  • A pace penalty: $(\dot{\gamma}_{i}-1)^{2}$, penalizes deviation from the desired pace.
  • A coordination penalty: $(\gamma_{i}-\gamma_{j})^{2}$ penalizes for the discrepancies with neighbors.
  • A control effort penalty: $(\ddot{\gamma}_{i})^{2}$ penalizes excessive control inputs.

### v2

@@ Abstract
Coordinated missions involving Unmanned Aerial Vehicles (UAVs) in dynamic environments pose significant challenges in maintaining both coordination and agility. In this paper, relying on the cooperative path following framework and using a game-theoretic formulation, we introduce a novel and scalable approach in which each UAV acts autonomously in different mission conditions. This formulation naturally accommodates heterogeneous and time-varying objectives across the system. In our setting, each UAV optimizes a cost function that incorporates temporal and mission-specific constraints. The optimization is performed within a one-dimensional domain, significantly reducing the computational cost and enabling real-time application to complex and dynamic scenarios. The framework is distributed in structure, enabling global, system-wide coordination (a Nash equilibrium) by using only local information. For ideal systems, we prove the existence and the Nash equilibrium exhibits exponential convergence. Furthermore, we invoke model predictive control (MPC) for non-ideal scenarios. In particular, we propose a discrete-time optimization approach that tackles path-following errors and communication failures, ensuring reliable and agile performance in dynamic and uncertain environments. Simulation results demonstrate the effectiveness and agility of the approach in ensuring successful mission execution across diverse realistic scenarios.
@@ I Introduction
Recent advancements in technology have significantly expanded the capabilities of Unmanned Aerial Vehicle (UAV) systems, enabling them to perform a wide range of tasks [27, 36]. Particularly, cooperative UAV systems have been extensively deployed across diverse domains, including civilian applications such as surveillance, environmental monitoring, and air traffic management, as well as military operations like swarm-based attack-defense scenarios. Notable examples include cooperative forest fire monitoring and suppression [29, 35], surveillance of multiple moving targets [14], to name a few. Despite their broad range of applications, cooperative UAV systems face operational limitations and challenges like battery endurance, payload carrying capability, flight autonomy, path planning, path following, and achieving reliable cooperation.
@@
Effective cooperation among a system of UAVs is crucial for accomplishing complex tasks that exceed the capabilities of a single UAV. Applications such as simultaneous target tracking [3], formation flying [23], and large-scale area mapping and monitoring [33] heavily rely on the coordinated efforts of multiple UAVs. Achieving reliable cooperation, however, is inherently challenging due to factors such as communication delays, limited computational resources, variations in UAV dynamics, disturbances, and unpredictable events. These challenges highlight the need for efficient, scalable and adaptive strategies to ensure system-wide collaboration in dynamic and uncertain environments. This paper addresses the challenge of time-critical cooperation by a novel game-theoretic method.
We develop a scalable distributed coordination framework for multi-UAV systems operating in realistic and dynamic environments. Since communication networks and disturbances are time-varying and uncertain, complete global information is generally unavailable. As a result, each UAV must retain local autonomy and adapt its objective online based on its own state and interactions with neighbors. This naturally motivates a game-theoretic formulation, which accommodates heterogeneous objectives, such as situations where some UAVs prioritize collision avoidance while others focus on temporal coordination (see Section IV-A4). Moreover, the proposed approach is inherently distributed and scalable: each UAV solves a low-dimensional local problem whose complexity does not increase with the size of the network, as neighbor information enters only through the cost function.
Following the framework of cooperative path-following developed in [2, 15, 16, 10], we formulate the time-critical coordination problem as a ”time synchronization” problem (see Section II for more details). This approach is inspired by the pioneering work of Leslie Lamport on time synchronization [21]. The key concept underpinning this framework is the decoupling of space and time in the general problem formulation, thereby reducing the multi-dimensional coordination problem to a one-dimensional consensus problem. This significantly reduces computational complexity and communication overhead, facilitating efficient real-time implementation. In [15, 18, 9, 16, 19] the consensus is achieved using a PI controller. The exponential stability of time coordination was established for networks connected in an integral sense. Despite the advantages of PI-based methods [16], they heavily rely on apriori mission planning, which weakens its adaptability and applicability in dynamic environments, particularly in satisfying operational requirements and time-varying objectives of the mission as they change. To overcome these limitations, in this paper, we reformulate the coordination problem within a game-theoretic framework, employing the virtual time concept (see [16]) as a consensus parameter. In our setting, UAVs engage strategically to reach agreement on the coordination parameter. Compared to previous PI-based methods [21], our approach provides greater generality and flexibility in accommodating operational constraints and dynamically evolving mission objectives. In particular, explicit constraints such as UAV speed and acceleration limits can be directly incorporated into the optimization problem, ensuring feasibility, unlike in PI-based methods, [16], where feasibility is more difficult to guarantee. Moreover, the optimization problem also enables the direct incorporation of complex, time-varying mission specifications—such as energy efficiency and collision avoidance (see Section III-B1)—into the optimization framework, further enhancing the adaptability and practicality of the proposed solution. Furthermore, our approach is distributed by its nature; hence, it maintains low per‐agent computational and communication overhead, ensuring scalable and robust coordination even in large‐scale, UAV networks (see Table II in Section III-B). Hence, this paper proposes a computationally efficient, scalable, and agile coordinated path-following method for UAV systems. A qualitative comparison between optimization-based and PI-based coordination methods is summarized in Table I.
TABLE I: Comparison between optimization-based (MPC) and PI-based coordination methods.
Feature Optimization-based PI-based
Path-following error Yes Yes
Explicit constraint handling Straightforward Depend on parameters choice
Operational requirements Yes No
Formal stability guarantees Guaranteed Depend on parameters choice
Computational requirement IPOPT solver None
Communication load Vector $\mathcal{N}_{i}\times K$ Vector $\mathcal{N}_{i}$
Ease of implementation
Figure 1 illustrates the architecture of the proposed solution. Unlike prior works [2, 15, 16] that employed a PI control law without addressing optimality, our approach introduces an Optimal Temporal Coordination block based on a game-theoretic formulation, where UAVs coordinate through constrained optimization, ensuring both feasibility and responsiveness to evolving mission demands. The main contributions of this paper are summarized as follows. We propose a novel game-based scalable coordination algorithm, where the Nash equilibrium of the strategic interactions among agents ensures system-wide coordination and successful task execution. We prove the existence of a Nash equilibrium and establish the exponential convergence of the solution under ideal conditions. We develop an MPC-based algorithm to handle realistic constraints, including path-following errors, communication failures, and dynamic environmental conditions.
In contrast to classical Nash equilibrium existence results (see, for example, [24], [28], and [39]), we establish the existence of a continuous-time, infinite-horizon Nash equilibrium in an infinite-dimensional action space with constraints. A related result—establishing the existence of a Nash equilibrium with bounded control inputs and stability—was previously shown in [38] for the discrete-time case using Lyapunov stability techniques. In this paper, we make a significant progress by proving both the existence and exponential stability of a continuous-time, infinite-horizon Nash equilibrium with constraints, based on the analysis of the associated Euler–Lagrange equations—an approach that, to the best of our knowledge, is used for the first time in this setting. In particular, our coordination problem leads to double integrator dynamics, resulting in a fourth-order system of Euler–Lagrange differential equations, which we study by deriving its explicit solution. Furthermore, by leveraging this explicit form and applying appropriate transversality conditions, we establish the exponential stability of the equilibrium.
Regarding Nash equilibrium seeking algorithms, there are existing results assuming a fixed communication network topology (see, for example [25, 34, 13, 37]). These methods often rely on standard consensus protocols or gradient-based dynamics in static environments, limiting their applicability in dynamic or time-varying settings. Recent works have begun to relax this assumption, considering time-varying network structures. For instance, the algorithms in [31, 5] explore convergence under dynamically evolving graphs, relying on conditions like joint or integral connectivity. However, these methods typically employ first-order dynamics and gradient-based updates. In this paper, leveraging MPC, we propose an algorithm (see Section III-B) that converges to Nash equilibrium in the case of time-varying network, even in the presence of communication failures. Furthermore, the algorithm takes into account the path following errors, making it robust to environmental disturbances such as wind (see Section IV for simulations). Nevertheless, each MPC step solves a simple optimization problem with the UAV’s virtual time as the only decision variable, avoiding optimization over high-dimensional UAV states and significantly reducing the computational cost. In the cooperative path-following framework, once the spatial trajectory is fixed during mission planning, the coordination layer operates solely on the virtual-time variable with simple upper and lower bounds, while neighbor information and path-following errors enter only as constants (see Problem 3). As a result, the method remains adaptive, robust, and scalable—its computational time is unaffected by the number of neighboring UAVs (see Table II)—making it suitable for real-time implementation. All computation times reported for each simulation scenario (see Section III-B) are on the order of $10^{-2}$, further evidencing the effectiveness of the proposed method.
@@
The remainder of this paper is structured as follows: Section II introduces the necessary notations and provides an overview of the preliminaries, including path planning, path following, and coordination. Furthermore, in Section II, we introduce an infinite horizon game problem with a discount factor. In Section III, by leveraging the system of Euler-Lagrange equations (a coupled fourth-order differential equations), we prove the existence of a Nash equilibrium for the corresponding unconstrained problem under ideal conditions. Moreover, we derive an explicit expression for the Nash equilibrium and prove that the equilibrium trajectory exhibits exponential convergence; see Proposition 3. Eventually, examining the coefficients of the explicit solution, we prove that there exists a discount rate such that the solution to the unconstrained problem is also solution to the constrained one; see Theorem 4. Section III-B outlines the general algorithm derived from the theoretical results, and Section IV validates the algorithm’s reliability and performance through simulations in challenging and realistic scenarios. The video footage showcasing the simulations and the flight experiments is available at the following link.
@@ II Preliminaries
We consider a class of time-critical cooperative missions that can be formulated as “time synchronization” problems (see, for example, [21, 16]). We assume that for a given mission a path planner generates a desired and feasible trajectory (possibly optimal for the mission) for each UAV. Next, we assume that each UAV is equipped with a robust low-level controller for tracking purposes. The coordination task is then reduced to a “time synchronization” problem. Although the proposed framework is general and can be applied to various heterogeneous UAV systems (multirotor, fixed-wing, VTOL, etc.), we focus on systems consisting of quadrotors for simplicity. Throughout the remainder of the paper, the term UAV refers to a quadrotor unless stated otherwise. Section II-A discusses the constraints associated with offline trajectory-generation, while Section II-B discusses constraints on coordination parameters and defines the objectives underlying the time-coordination formulation.
Figure 1: Game-theoretic cooperative path-following approach block diagram.
@@ II-A Path Planning and Following
Consider a system of $N$ UAVs involved in the mission, $N\in\mathbb{N}$. For each UAV, the desired trajectory is given as a function $x_{d,i}:[0,t_{d,i}^{*}]\rightarrow\mathbb{R}^{3}$, $t_{d,i}^{*}>0$. For all $t\in[0,t_{d,i}^{*}]$, the trajectories satisfy the constraints defined by the $i^{\rm th}$ UAV’s minimum and maximum linear velocity limits, and maximum linear acceleration limits [16, 8]:
[…]
@@
[…]
@@
Time-critical mission specifications are encoded into the desired trajectories of the UAVs by the planner. By properly defining these trajectories, it is possible to design missions with varying characteristics, such as simultaneous arrival and/or sequential auto-landing. The analysis of such systems is similar, and hence, here, we consider only the simultaneous arrival case. All UAVs are expected to complete their assigned paths under ideal conditions sharing a common nominal predefined final arrival time $t_{d}^{*}\in\mathbb{R}^{+}$: $t_{d,i}^{*}=t_{d}^{*}$, $i=1,\ldots,N.$ Next, we assume that the trajectories maintain spatial separation throughout the mission:
[…]
where $E>0$ denotes the prescribed minimum allowable distance between any two UAV trajectories.
Remark 1. In Section III-B1, we introduce an additional collision-avoidance term within the optimization framework. This term explicitly enforces safe separation between UAVs, thereby eliminating the need for the spatial separation constraints in (3). Consequently, the proposed formulation allows UAVs to follow intersecting or closely aligned trajectories while ensuring collision-free operation.
@@ II-B Time Coordination for Simultaneous Arrival
To formulate the simultaneous arrival problem, we follow [9, 16]. Let $\gamma_{i}:\mathbb{R}^{+}\rightarrow[0,t_{d}^{*}]$, $i=1,\dots,N$, map the clock time $t$ to the mission time $t_{d,i}=\gamma_{i}(t)$, where $t_{d,i}$ parameterizes a predefined desired trajectory. The function $\gamma_{i}$ is referred to as the virtual time of the $i$th UAV and represents an internal coordination variable that governs the progression of the UAV along its path. Unlike real time, $\gamma_{i}$ can be adjusted dynamically, enabling reparameterization of the desired trajectory as
$x_{\gamma,i}(t):=x_{d,i}(\gamma_{i}(t)),$ (4)
where $x_{\gamma,i}$ denotes the virtual target to be tracked by the UAV. This formulation decouples temporal evolution from spatial motion, allowing coordination through appropriate regulation of the virtual time. Specifically, When $\dot{\gamma}_{i}(t)=1$, the UAV travels at the desired pace. When $\dot{\gamma}_{i}(t)>1$, the UAV moves faster than the desired pace. When $\dot{\gamma}_{i}(t)<1$, the UAV moves slower than the desired pace. Thus, $\gamma_{i}$ serves the role of a consensus parameter in our problem. Based on the physical limitations and mission requirements of the UAVs, as defined in (1) and (2), we derive general bounds for the consensus parameter of each UAV. Specifically, from (4), we have
@@
The first equation along with (1) implies that the minimum and the maximum values of all admissible parameters $\gamma_{i}$ must satisfy the following constraint (for more details see [9]):
@@
Note that since $v^{i}_{d,\max}<v^{i}_{\max}$ and $a^{i}_{d,\max}<a^{i}_{\max}$, there exist $\gamma_{i}$ parameters such that inequalities in (6) and (7) are satisfied. For more details on above derivation see [9].
@@ II-C Problem Formulation
Next, we introduce a game-theoretic approach that facilitates time-critical cooperative mission execution. For the theoretical analysis, we make the following assumptions regarding the UAV system and its dynamics:
Assumption 1
Communication network of UAVs is fully connected; that is, $\mathcal{N}_{i}=\{1,\ldots,N\}$ for all $i=1,\ldots,N$, where $\mathcal{N}_{i}$ is set of UAVs with which $i^{th}$ UAV can exchange information.
Assumption 2
The motion of each UAV is governed by ideal path-following kinematics, meaning that the UAV perfectly tracks its assigned geometric path without deviation.
Remark 1
Assumption 2 implies that the desired trajectory reparameterizations through the corresponding virtual times $\gamma_{i}$, together with the linear constraints in (6) and (7), still yield feasible trajectories. Hence, no path-following error occurs under this assumption.
While these assumptions simplify the theoretical analysis (see Section III-A), they do not limit the practical applicability of the proposed approach. In Section III-B, we introduce an algorithm that incorporates path-following errors and communication failures, enabling application in real-world scenarios. As shown in [7, 17], by using cascaded inner-loop outer-loop structure for flight control applications, the uncertainties in system dynamics can be handled by a variety of robust control methods.
We begin by introducing the admissible set of the consensus parameter (virtual time) for each UAV. Due to unexpected events, UAVs may begin their missions at different time instants, which corresponds to different initial virtual times, i.e., $\gamma_{i}(0)=\gamma_{i}^{0}$ for $i=1,\dots,N$. Along with the inequalities in (6), (7), we define the admissible sets:
$\mathcal{A}^{0}_{i}:=\Big\{\gamma_{i}\in H^{2}_{w,\alpha}((0,\infty)):\gamma_{i}(0)=\gamma^{0}_{i}\geqslant 0,\,\dot{\gamma}_{i}(0)=1\Big\},$
@@
where $H^{1}_{w,\alpha}((0,\infty))=\{g\in H^{2}_{loc}((0,\infty)):\int_{0}^{\infty}e^{-\alpha t}(g^{2}+\dot{g}^{2}+\ddot{g}^{2}){\rm d}t<\infty\}$ is weighted Sobolev space (for a detailed discussion of Sobolev and weighted Sobolev spaces see for example [1], [20] and [30]) with weight $e^{-\alpha t}$.
@@
$\begin{split}J_{i}(\gamma_{i},\gamma_{-i})=\int_{0}^{\infty}&e^{-\alpha t}\Big(w_{1}\left(\dot{\gamma}_{i}-1\right)^{2}\\ &+\frac{w_{2}}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}(\gamma_{i}-\gamma_{j})^{2}+w_{3}\ddot{\gamma}_{i}^{2}\Big){\rm d}t,\end{split}$ (10)
where $\gamma_{-i}=(\gamma_{1},\dots,\gamma_{i-1},\gamma_{i+1},\dots,\gamma_{N})$, $\mathcal{N}_{i}$ is set of neighbors of $i^{th}$ UAV and $w_{1},w_{2},w_{3}\in\mathbb{R}_{0}^{+}$ are nonnegative constants satisfying $w_{1}+w_{2}+w_{3}=1$, which determine the relative importance of the corresponding penalty terms.
The cost function in (10) corresponds to simultaneous arrival and maintaining a desired mission pace with the following terms. Pace penalty: $(\dot{\gamma}_{i}-1)^{2}$ penalizes deviations from the desired pace, and the parameter $w_{1}$ determines its relative importance. Coordination penalty: $(\gamma_{i}-\gamma_{j})^{2}$ penalizes discrepancies with neighboring agents, and the parameter $w_{2}$ determines its relative importance. Control effort penalty: $(\ddot{\gamma}_{i})^{2}$ penalizes excessive control inputs, and the parameter $w_{3}$ determines its relative importance.